OpenStudy (anonymous):

A Customer went to a garden shop and bought some potting soil for $12.50 and 6 shrubs. If the total bill was $71.00, how much did each shrub cost if they all cost the same. Choose a label a variable for this situation. Write a equation using the variable that would answer the question. Solve your equation and check your answer. Use the solution to write a sentence answer to the word problem/.

OpenStudy (lilg132):

Total bill - potting soil = X x = prices of 6 shrubs x / 6 = price of each shrub dont forget to click good answer

OpenStudy (anonymous):

what do i divide?

OpenStudy (lilg132):

ok so have you worked out what x is yet? to work out x you do Total bill - potting soil price

OpenStudy (anonymous):

So divide 71 divided by 6?

OpenStudy (anonymous):

Im mixed up

OpenStudy (lilg132):

no wait your rushing ahead

OpenStudy (lilg132):

what i said was the first step to do is TOTAL BILL - THE PRICE OF POTTING SOIL

OpenStudy (lilg132):

so what was the total bill?

OpenStudy (lilg132):

and what was the price of potting soil?

OpenStudy (anonymous):

the bill was $71.00 and the soil was $12.50

OpenStudy (lilg132):

great so all you do is total bill - price of potting soil which is the same as 71.00 - 12.50 = ??

OpenStudy (anonymous):

58.50

OpenStudy (anonymous):

It is 58.50??

OpenStudy (lilg132):

yes

OpenStudy (lilg132):

now that means we have worked out how much 6 shrubs cost

OpenStudy (anonymous):

it costs 58.50?

OpenStudy (lilg132):

yes but to work out the price of 1 shrub what do you think we do if that was the price of 6 shrubs

OpenStudy (lilg132):

6 shrubs = 58.50 1 shrub = 58.50 / 6

OpenStudy (anonymous):

$9.75

OpenStudy (lilg132):

great

OpenStudy (anonymous):

So

OpenStudy (lilg132):

ok now to write an equation we just replace the answer that we dont know by x so for example the first step we done which was Total bill - price of potting soil = ??? because we dont know the answer we can call it x

OpenStudy (lilg132):

do you get that so far?

OpenStudy (anonymous):

71 + 12.50 is 83.50

OpenStudy (lilg132):

but obviously we do know the answer because we already worked it out but lets just pretend we are seeing this question for the first time

OpenStudy (lilg132):

its actually 71 - 12.50

OpenStudy (anonymous):

So the answer would be 71.00 - 12.50 = 58.50

OpenStudy (lilg132):

we're not after the answer at this stage but yes replace 58.5 with x

OpenStudy (lilg132):

so 71 - 12.5 = x

OpenStudy (anonymous):

X = 58.50

OpenStudy (lilg132):

we are trying to work out the equation now not the answer

OpenStudy (lilg132):

thats right x = 58.50

OpenStudy (lilg132):

now the equation we can write is ps = per shrub price x = 71 - 12.5 x / 6 = PS so now that is the equation we will use x / 6 = ps

OpenStudy (lilg132):

and to solve using that equation we need to know what x is first remember x = total bill - price of potting soil which is the same as x = 71 - 12.5 you already worked out x and you got x = 58.5

OpenStudy (lilg132):

so now in the equation where it says x just replace it by the number 58.5 x / 6 = PS which will look like 58.5 / 6 = PS now to get the answer for PS you just calculate what is written which is 58.5 divided by 6 will give you PS

OpenStudy (anonymous):

You there?

OpenStudy (lilg132):

yes

OpenStudy (anonymous):

You there now?

OpenStudy (anonymous):

Sorry was busy